What is Z score normalization?
The z Score utility model transforms the distribution of pixel values into a standard normal distribution (z-score value). By this normalization the images of different individuals become more comparable.
For more information on z-score, check this article:
http://en.wikipedia.org/wiki/Standard_score
What is the probability of rolling a composite number with one roll of a number cube?
The composite numbers on an ordinary cube are 4 and 6 so the probability of rolling a composite is 2/6 = 1/3
A single die is toss Find the probability of getting a 5?
A 6-sided die has 6 possible outcomes. If each of the sides has a different number from 1 to 6, then the probability of getting any one number is 1/6 or 16.67%.
P(A) = 4/52 or 1/13 or 0.077 and P(B) = 13/52 or 1/4 or .25.
You must remove 12 socks to be sure of having a pair of Brown one
How do you check for normal distribution?
There are two main methods: theoretical and empirical.
Theoretical: Is the random variable the sum (or mean) of a large number of imdependent, identically distributed variables? If so, by the Central Limit Theorem the variable in question is approximately normally distributed.
Empirical: there are various goodness-of-fit tests. Two of the better known are the chi-square and the Kolmogorov-Smirnov tests. There are others. These compare the observed values with what might be expected if the distribution were Normal. The greater the discrepancy, the less likely it is that the distribution is Normal, the smaller the discrepancy the more likely that the distribution is Normal.
Can a probability be a number greater than 1?
no, because lets say 1 equals 100% if there is a 1.5(150%) chance of sun today it means its happening no matter what but also 1 means it is guarantee it will be to so really it doesn't make a difference
The answer depends on how many cards are in the hand.
It is 1/18.
The answer depends on what the trial comprises. Since that is not defined, there can be no sensible answer.
Since no two notes can sum to another, the answer is 25 - 1 = 31.
What is the secret to calculating probabilities of compound events?
There is no secret: the procedures are well studied. However, it is important to know whether the events are independent or dependent.
50% or 3 out of 6.
Let's say there are 52 cards in a deck
There's one chance out of 52 to get the two of clubs from the first deck
The same chance with the second deck
The probability that the two events occur at the same time is the probability of the first multiplied by the probability of the second as the events are independent
So P = 1/52 x 1/52 = 1/2074
The probability is 19/25 * 18/24 = 0.57
What is the probability of getting three tails and a head when a coin is tossed four times?
The probability is 1/4
8 in 12 for the greater than 3 case. (4, 5, 6, 7, 8, 9, 10, 11, 12)
6 in 12 for the even case. (2, 4, 6, 8, 10, 12)
9 in 12 for the combined case. (2, 4, 5, 6, 7, 8, 9, 10, 11, 12)
If the sample size is less then 30 use the T table, if greater then 30 use the Z table.
A Leap year has 366 days. in which you have 52 weeks and 2 days.
the 2 days may be
sun,Mon
mon,Tue
tue,wed
wed,THu,
thu,Fri
FRi,SAT
sat,sun
so you have 7 options among which 2 u can choose..
so the answer is 2/7 for having 53 Sundays. The probability of having 53 Thursdays is also 2/7.
The probability of having either 53 Sundays or 53 Thursdays is 4/7.
The probability of getting a spade out of a full deck is 13/52 = 1/4.
Now, to get exactly one spade you have the following possibilities:
a. Spade first, then others: 13/52*39/51*38/50
b. Other, then Spade, then other: 39/52*13/51*38/50
c. Other, then other, then spade: 39/52*38/51*13/50.
Sum them up, and you get 3*13*39*38/(52*51*50) = 0.4385
To understand this best, please draw a tree.
Relation between mean and standard deviation?
Standard deviation is the variance from the mean of the data.
How many black 2's and red 3's in a deck of 52 cards?
There are 2 black 2's and 2 red 3's in a standard deck of 52 cards.